L^{A}TEX News

Issue 7, June 1997

T1 encoded Computer Modern fonts

As in the last release the base L^{A}TEX distribution contains three different sets of ‘fd’ files for T1 encoded fonts.

In this release the default installation uses ec.ins and so installs files suitable for the current ‘EC fonts’ distribution. If you have still not updated to the EC fonts and are using the earlier test versions, known as DC then you should unpack newdc.ins (for DC release 1.2 or later) or olddc.ins (for the original releases of the DC fonts). This should be done after unpacking unpack.ins but before making the format by running iniTEX on latex.ltx. There are further details in

install.txt.

T1 encoded Concrete fonts

The Metafont sources for T1 encoded ‘Concrete’ fonts have been removed from the mfnss distribution as they were based on the now obsolete DC fonts release 1.1. Similarly the cmextra.ins install file in the base distribution no longer generates fd files for the ‘Concrete’ fonts. To use these fonts in either T1 or OT1 encoding it is recommended that you obtain Walter Schmidt’s ccfonts package and fonts from CTAN

macros/latex/contrib/supported/ccfonts.

Further input encodings

Two more inputenc packages have been added: for latin5, thanks to H. Turgut Uyar; and for latin3, thanks to Jörg Knappen.

Normalising spacing after punctuation

The command \normalsfcodes was introduced at the last patch release. This is normally given the correct definition automatically and so need not be explicitly set. It is used to correct a problem, reported by Donald Arseneau, that punctuation in page headers has always (in all known TEX formats) been potentially incorrect if the page break happens while a local setting of the space codes (for instance by the command \frenchspacing) is in effect. A common example of this happening in L^{A}TEX is in the verbatim environment.

Accessing Bold Math Symbols

The tools distribution contains a new package, bm, which defines a command \bm that allows individual bold symbols to be accessed within a math expression (in contrast to \boldmath which makes whole math expressions default to bold fonts). It is more general than the existing amsbsy package; however, to ease the translation of documents between these two packages, bm makes \boldsymbol an alias for \bm.

This package was previously made available from the ‘contrib’ area of the CTAN archives, and as part of Y&Y’s L^{A}TEX support for the MathTime fonts.

Policy on standard classes

Many of the problem reports we receive concerning the standard classes are not concerned with bugs but are suggesting, more or less politely, that the design decisions embodied in them are ‘not optimal’ and asking us to modify them.

There are several reasons why we have decided not to make such changes to these files.

• However misguided, the current behaviour is clearly what was intended when these classes were designed.

• It is not good practice to change such aspects of ‘standard classes’ because many people will be relying on them.

We have therefore decided not to even consider making such modifications, nor to spend time justifying that decision. This does not mean that we do not agree that there are many deficiencies in the design of these classes, but we have many tasks with higher priority than continually explaining why the standard classes for L^{A}TEX cannot be changed.

We would, of course, welcome the production of better classes, or of packages that can be used to enhance these classes.

New addresses for TUG

For information about joining the TEX Users Group, and about lots of other L^{A}TEX-related matters, please contact them at their new address:

TEX Users Group, P.O. Box 1239, Three Rivers, CA 93271-1239, USA Fax: +1 209 561 4584

                                                             E-mail: tug@mail.tug.org
                                                             URL: http://www.tug.org/

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 1997, all rights reserved.